OpenStudy (anonymous):

Solve for A; 2 times the square root of (a-1)=4 times the square root of a

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt{a-1}=\frac{4\sqrt{a}}{2}\] \[\sqrt{a-1}=2\sqrt{a}\] squaring on both sides \[(\sqrt{a-1})^2=(2\sqrt{a})^2\] \[a-1=4a\] now solve for a

OpenStudy (anonymous):

ok thanks